# LLM-Powered Workflow Optimization for Multidisciplinary Software Development: An Automotive Industry Case Study

### 22 Mar 2026 v1

@@ 1. Introduction
In many industrial software projects, domain experts and software developers must collaborate across disciplinary boundaries, coordinating through heterogeneous artifacts that each party produces, maintains, and evolves independently. Experts contribute specifications, signal definitions, and design documents grounded in their respective domains, while developers must reconcile these artifacts and translate them into a unified, executable codebase. This coordination pattern is straightforward when artifacts are few and stable, but becomes a persistent bottleneck as systems grow in scope and the number of contributing disciplines increases. The automotive industry, with its deep regulatory requirements, safety-critical constraints, and multidisciplinary teams, offers an instructive example of how document-to-code translation can dominate engineering effort (Wang et al., 2025a).
@@
This pattern is characteristic of Multidisciplinary Software Development (MSD), which arises whenever software systems must encode specialized knowledge from non-software domains. In MSD settings, domain experts (engineers, scientists, regulators, or business analysts) produce specifications using discipline-specific formalisms that software developers must interpret, align, and implement as coherent system behavior (Li et al., 2017; Mougin et al., 2015). The collaboration is inherently two-way: developers translate domain requirements into code, while domain experts must understand implementation constraints to refine their specifications (Zheng et al., 2018; Guerineau et al., 2017). This continuous dialogue creates value by embedding deep expertise into software, but it also introduces persistent operational challenges that scale poorly with system complexity.
@@
Figure 1 illustrates a common pattern we observe in MSD: requirements are authored across heterogeneous documents by different stakeholders, then manually translated by developers into implemented endpoints, followed by repeated feedback cycles where feasibility and constraints are communicated back to domain experts for refinement. In spapi, this translation spans a patchwork of documents and a correspondingly tedious patchwork of implementation tasks, which slows delivery and increases the risk of defects. Our goal, shown in Figure 1(b), is to automate the translation from domain artifacts to spapi endpoints. Achieving this requires transforming the MSD workflow itself: we reorganize the handoffs and dependencies so that LLM-powered services can replace repeated manual coordination. We design this workflow transformation using a graph-based representation of the artifacts and their dependencies.
We make four contributions relevant to practitioners facing similar MSD challenges.
@@
The remainder of this paper is organized as follows. Section 2 describes the spapi system and the operational problems observed in its development workflow. Section 3 presents our workflow optimization approach and its implementation. Section 4 reports experimental results and stakeholder feedback. Section 5 discusses practical considerations. Section 6 surveys related work, and Section 7 concludes.
@@ 2. The spapi System
spapi is an in-vehicle web server deployed at Volvo Group that provides a stable, authenticated API layer mapping vehicle signals and control state to well-typed RESTful endpoints. Driver-facing mobile applications and backend fleet services consume these endpoints to display vehicle status, adjust climate and comfort settings, and monitor operational parameters across the vehicle’s functional domains.
Developing and maintaining spapi has required 15 to 20 full-time engineers (FTEs) to deliver over 100 APIs spanning six functional domains: driver productivity, connected systems, energy management, vehicle systems, visibility, and dynamics. Three categories of artifacts drive this workflow. OpenAPI specifications define endpoint behavior, data types, and interface contracts; these artifacts embody the essential complexity (Jr., 1987) of the system and serve as the authoritative source for mobile and cloud application integration. Signal definitions describe low-level CAN bus messages encoding vehicle state, authored by control engineers and system architects who understand the physical systems. Mapping documents bind high-level API properties to their underlying signals, specifying how abstractions like “climate mode” translate to specific CAN message fields.
@@
Manual-translation overload. Developers spend substantial effort transcribing information from specifications and signal definitions into code, diverting time from higher-value activities such as design, testing, and verification. Even with assistance from coding agents like GitHub Copilot, developers must still manually locate relevant artifacts, interpret domain-specific formalisms, reconcile inconsistencies across documents, and repeatedly consult domain experts to resolve ambiguities. This coordination burden grows proportionally with API and signal count. With over 400 unique properties and 776 associated CAN signals across the spapi system, translation work creates a persistent bottleneck that cannot be resolved by coding assistants alone; it requires automating the workflow itself.
Ambiguity-driven churn. Underspecified items in domain documents prompt clarification rounds between developers and experts, each consuming time from both parties and potentially invalidating prior implementation decisions. Signal definitions in spapi were often concise to the point of ambiguity, requiring developers to consult vehicle state experts, system engineers, and sometimes control engineers, whose discipline has deep technical foundations spanning decades (Bennett, 1993). A single ambiguous signal could cascade into multiple days of clarification before implementation could proceed.
Coordination bottlenecks. The coordination group established to streamline information flow across spapi development proved inadequate for managing hundreds of API properties and vehicle signals. Developers began bypassing the group to contact domain experts directly, creating ad-hoc communication patterns that were difficult to track and prone to inconsistency. Once such workarounds begin delivering value, organizational inertia makes them resistant to change, even when their inefficiency is widely recognized.
@@ 3.4.2. Signal-Property Synthesis
The second service aligns API properties with their corresponding CAN signal handlers. A key challenge in synthesizing deployable API endpoints lies in correctly aligning API properties with underlying CAN signals. Given an API property defined at the application level, the system must identify one or more corresponding CAN signals and determine how their values should be interpreted, transformed, and composed to produce semantically correct API behavior.
This alignment is challenging because API and CAN specifications often diverge. Value domains may not match (e.g., ON/OFF vs. 0/1), units may differ (e.g., m/s on CAN vs. km/h in the API), and some properties aggregate multiple CAN signals (e.g., a single time property derived from separate minute and second signals). As a result, naive matching based on names or types is unreliable and can introduce silent semantic errors.
To make these assumptions explicit and verifiable, we use a JSON alignment template as an intermediate representation. The LLM must enumerate all contributing CAN signals and specify explicit value mappings, enum correspondences, and unit conversion annotations when needed. As shown in Figure 4, we further guide generation with a chain-of-thought prompt that encourages stepwise reasoning over (i) property intent, (ii) value and unit consistency, and (iii) whether the mapping is direct, requires transformation, or involves multi-signal composition. Together, the template constraint and structured reasoning reduce hallucinated assumptions and improve robustness across heterogeneous signal specifications.
@@ 3.4.3. Property-Endpoint Synthesis
The third service assembles property handlers into deployable API endpoints based on the OpenAPI specification. Given the API contract and generated property accessors as input, it produces complete FastAPI endpoint implementations, including routing, request validation, response serialization, and error handling.
Synthesizing endpoints directly from specifications is error-prone in practice. Naively generated endpoints may violate the OpenAPI contract, for example by using inconsistent parameter names, omitting or introducing response fields, or mismatching route paths and HTTP methods. In addition, validation logic such as range checks or enum constraints may be partially implemented or entirely omitted, leading to subtle contract violations that are difficult to detect at runtime.
To mitigate these risks, this service adopts a contract-first, template-based generation strategy. A FastAPI endpoint boilerplate template (Figure 5) fixes the endpoint structure, routing, schemas, and cross-cutting concerns such as authentication, logging, and error handling. The LLM is restricted to instantiating endpoint-specific logic within predefined slots, guided by the prompt shown in Figure 6. Validation is delegated to FastAPI’s schema-driven mechanisms, ensuring consistent enforcement of constraints specified in the OpenAPI document.
@@ 3.6. Deployed System
After iterative optimization, the final system replaces all manual translation steps with automated services and reaches a stable configuration validated by domain experts. The resulting workflow, shown in the lower-right of Figure 2, consists of three server nodes connected in a linear pipeline. Compared to the initial manual workflow, the deployed system achieves a net complexity reduction of 5, a communication quality improvement of 6, and an automation increase of 3.
@@
The system takes as input an initial workflow graph $\mathcal{G}^{(0)}$ with dense manual coordination and a set of domain artifacts, including OpenAPI specifications, signal definitions, and constraint documents. It outputs an optimized workflow graph $\mathcal{G}^{*}$ in which manual translation nodes are replaced by service nodes and redundant coordination edges are removed, yielding a compact and executable end-to-end pipeline.
@@
Each node in $\mathcal{G}^{*}$ is deployed as an independent server exposing a well-defined API and composed sequentially to mirror the optimized workflow graph. The architecture decouples workflow logic from the underlying language model implementation, enabling flexible substitution of open-source or proprietary LLMs under different cost, latency, and privacy constraints. A unified human–machine interface allows domain experts to inspect API-to-signal mappings, review generated code, and request regeneration with additional constraints, preserving human oversight without manual translation.
@@
The system supports both fully automated and human-in-the-loop execution. For well-specified tasks, the pipeline operates end-to-end without intervention; ambiguous mappings or specification inconsistencies are flagged for expert review. By replacing input artifacts or domain-specific instructions, the same framework can be readily adapted to new APIs, vehicle platforms, or other translation-heavy workflows.
@@ 4.1. Experimental Setup
The baseline implementations have been reviewed and deployed in production and therefore reflect accepted system behavior. In this domain, semantic errors typically manifest as concrete issues such as enum mismatches, value range violations, or missing signal bindings, all of which are explicitly targeted by our validation and automated debugging mechanisms. We deliberately adopt a conservative matching strategy, preferring to flag ambiguous cases for human review rather than generating potentially incorrect code, in order to avoid silent semantic errors.
To address RQ4, we invited all practitioners with direct experience of both workflows to provide structured feedback: four domain experts and two developers. This complete census of qualified evaluators rated the system on role-specific criteria using a 5-point scale. Experts assessed communication effort, functional coverage, and implementation accuracy; developers assessed communication efficiency, debugging effort, and code maintainability.
@@
Since no existing automated solutions target API development in industrial automotive scenarios, we compare our system with APIs developed by human engineers assisted by proprietary AI coding coding agents. In this baseline, engineers were permitted to use GitHub Copilot, a state-of-the-art AI-assisted coding tool, to support implementation. We assess both API quality and development time.
@@
The system distinguishes between offline processing and deployment-time execution to manage LLM usage efficiently. All compute-intensive LLM operations occur during the offline stage and do not affect runtime performance. In our experiments, we use the proprietary LLM GPT-4o (2024-05-13).
During offline processing, the LLM is used to scan and filter the entire CAN database and derive read/write equations for relevant signals. In a full experimental run involving 192 APIs and 776 CAN signals, the system made approximately 15,000 LLM calls over 20.6 hours. This corresponds to an average cost of 396 seconds and 80 LLM calls per API.
During deployment, the previously generated read/write equations are reused and do not require regeneration. The LLM is only invoked for lightweight tasks such as matching API properties to signals and synthesizing the final API endpoint. On average, each API requires about 5 LLM calls, resulting in approximately 10 seconds per API. This overhead is low and acceptable for practical use in engineering workflows.
4.2. RQ1: Code Quality Across Domains
Table 3 summarizes code quality results across all six automotive domains. The automated workflow achieves strong performance consistently: four domains (Driver productivity, Energy, Visibility, and Dynamics) reach perfect precision of 100%, indicating that all generated property-level code was correct. The remaining two domains (Connected systems and Vehicle system) achieve precision above 94%.
Recall scores are slightly lower, ranging from 86.5% to 95.7% across domains. This reflects our conservative matching strategy during property-to-signal alignment, where we apply strict filtering to retain only high-confidence matches. Properties with ambiguous signal mappings are flagged for human review rather than generating potentially incorrect code. Despite this conservative approach, average recall exceeds 90%, indicating strong coverage of the original specifications.
The overall F1 score of 93.7% demonstrates that the automated workflow produces code quality comparable to baseline development across diverse automotive domains. Importantly, performance remains consistent regardless of which expert roles are involved–the system handles APIs requiring input from control engineers (Dynamics, Energy) as effectively as those involving HMI/UX designers or business translators.
Stakeholder satisfaction scores (detailed in Section 4.5) provide additional validation, with both experts and developers reporting high satisfaction and noting that the system facilitates transparent monitoring through well-defined interfaces.
@@ 5. Discussion
Although Section 3 provides technical criteria for workflow automation, selecting suitable workflows ultimately requires informed judgment. Beyond technical feasibility, successful adoption depends on organizational readiness and openness to change, which are often difficult to assess in advance. In the case of spapi, the workflow was both well-scoped and supported by a team willing to experiment, which proved critical to achieving tangible benefits.
@@
Our results show that LLMs can support multiple stages of workflow automation, including parsing, translation, generation, and evaluation. However, effective use of LLMs still requires experience, particularly in deciding when hybrid solutions combining scripts and LLMs are more appropriate. The iterative nature of our approach allows LLM usage to evolve alongside the workflow as teams refine these decisions over time.
@@ 7. Conclusion
In this paper, we show that LLM-powered workflow automation works in production MSD settings. By modeling workflows as graphs and selectively automating translation steps, we substantially reduce coordination overhead while preserving human oversight. The spapi deployment at Volvo Group validates this approach at scale. The pattern we address is not automotive-specific; it applies wherever software encodes specialized domain knowledge.

### 24 Mar 2026 v2

@@ 1. Introduction
In many industrial software projects, domain experts and software developers must collaborate across disciplinary boundaries, coordinating through heterogeneous artifacts that each party produces, maintains, and evolves independently. Experts contribute specifications, signal definitions, and design documents grounded in their respective domains, while developers must reconcile these artifacts and translate them into a unified, executable codebase. This coordination pattern is straightforward when artifacts are few and stable, but becomes a persistent bottleneck as systems grow in scope and the number of contributing disciplines increases. The automotive industry, with its deep regulatory requirements, safety-critical constraints, and multidisciplinary teams, offers an instructive example of how document-to-code translation can dominate engineering effort.
@@
This pattern is characteristic of Multidisciplinary Software Development (MSD), which arises whenever software systems must encode specialized knowledge from non-software domains. In MSD settings, domain experts (engineers, scientists, regulators, or business analysts) produce specifications using discipline-specific formalisms that software developers must interpret, align, and implement as coherent system behavior (Li et al., 2017). The collaboration is inherently two-way: developers translate domain requirements into code, while domain experts must understand implementation constraints to refine their specifications (Zheng et al., 2018). This continuous dialogue creates value by embedding deep expertise into software, but it also introduces persistent operational challenges that scale poorly with system complexity.
@@
Figure 1 illustrates a common pattern we observe in MSD: requirements are authored across heterogeneous documents by different stakeholders, then manually translated by developers into implemented endpoints, followed by repeated feedback cycles where feasibility and constraints are communicated back to domain experts for refinement. In spapi, this translation spans a patchwork of documents and a correspondingly tedious patchwork of implementation tasks, which slows delivery and increases the risk of defects. Our goal, shown in Figure 1(b), is to automate the translation from domain artifacts to spapi endpoints. Achieving this requires transforming the MSD workflow itself: we reorganize the handoffs and dependencies so that LLM-powered services can replace repeated manual coordination. We design this workflow transformation using a graph-based representation of the artifacts and their dependencies. We make four contributions relevant to practitioners facing similar MSD challenges:
@@
The rest of this paper is organized as follows. Section 2 describes the spapi system and its workflow challenges. Section 3 presents our optimization approach. Section 4 reports results and stakeholder feedback. Section 5 discusses practical considerations, Section 6 reviews related work, and Section 7 concludes.
@@ 2. The spapi System
spapi is an in-vehicle web server deployed at Volvo Group that provides an authenticated API layer mapping vehicle signals and control state to well-typed RESTful endpoints. Driver-facing mobile applications and backend fleet services consume these endpoints to display vehicle status, adjust comfort settings, and monitor operational parameters across the vehicle’s functional domains.
Developing and maintaining spapi has required 15 to 20 full-time engineers (FTEs) to deliver more than 100 APIs across six functional domains: driver productivity, connected systems, energy management, vehicle systems, visibility, and dynamics. Three types of artifacts drive this workflow. OpenAPI specifications define endpoint behavior, data types, and interface contracts; they capture the essential complexity of the system and serve as the authoritative source for mobile and cloud integration. Signal definitions describe low-level CAN bus messages that encode vehicle state, authored by control engineers and system architects with knowledge of the physical systems. Mapping documents bind high-level API properties to their underlying signals, specifying how abstractions like “climate mode” translate to specific CAN message fields.
@@
Manual-translation overload. Developers spend substantial effort transcribing information from specifications and signal definitions into code, drawing time away from higher-value work such as design, testing, and verification. Even with coding agents like GitHub Copilot, they must still locate relevant artifacts, interpret domain-specific formalisms, reconcile inconsistencies, and consult domain experts to resolve ambiguities. This burden scales with API and signal count. With over 400 unique properties and 776 associated CAN signals in the spapi system, translation remains a persistent bottleneck that coding assistants alone cannot remove; the workflow itself must be automated.
Ambiguity-driven churn. Underspecified domain documents trigger repeated clarification between developers and experts, consuming time and sometimes invalidating earlier implementation decisions. Signal definitions in spapi were often so concise as to be ambiguous, requiring consultation with vehicle state experts, system engineers, and sometimes control engineers, whose discipline has deep technical foundations spanning decades. A single ambiguous signal could delay implementation by days.
Coordination bottlenecks. The coordination group created to streamline information flow in spapi development proved insufficient for managing hundreds of API properties and vehicle signals. Developers began bypassing it to contact domain experts directly, creating ad hoc communication patterns that were hard to track and prone to inconsistency. Once such workarounds start delivering value, organizational inertia makes them difficult to replace, even when their inefficiency is widely recognized.
@@ 3.4.2. Signal-Property Synthesis
The second service aligns API properties with their corresponding CAN signal handlers. A key challenge in synthesizing deployable API endpoints is correctly aligning application-level API properties with underlying CAN signals. Given an API property, the system must identify the relevant CAN signals and determine how their values should be interpreted and composed to produce semantically correct behavior.
This is difficult because API and CAN specifications often diverge. Value domains may differ (e.g., ON/OFF vs. 0/1), units may not match (e.g., m/s on CAN vs. km/h in the API), and some properties aggregate multiple CAN signals (e.g., a time property derived from separate minute and second signals). So, naive matching by names or types is unreliable and may introduce silent semantic errors.
To make these assumptions explicit and verifiable, we use a JSON alignment template as an intermediate representation. The LLM must enumerate all contributing CAN signals and specify explicit value mappings, enum correspondences, and unit conversions when needed. As shown in Figure 4, generation is further guided by a chain-of-thought prompt encouraging stepwise reasoning over (i) property intent, (ii) value and unit consistency, and (iii) whether the mapping is direct, transformed, or composed from multiple signals. Together, the template constraint and structured reasoning reduce hallucinated assumptions and improve robustness across heterogeneous signal specifications.
@@ 3.4.3. Property-Endpoint Synthesis
The third service assembles property handlers into deployable API endpoints based on the OpenAPI specification. Given the API contract and generated property accessors, it produces complete FastAPI endpoint implementations, including routing, request validation, response serialization, and error handling.
Synthesizing endpoints directly from specifications is error-prone. Naively generated endpoints may violate the OpenAPI contract by using inconsistent parameter names, omitting or introducing response fields, or mismatching route paths and HTTP methods. Validation logic such as range checks or enum constraints may also be partially implemented or omitted, leading to subtle contract violations that are hard to detect at runtime.
To mitigate these risks, this service adopts a contract-first, template-based generation strategy. A FastAPI endpoint boilerplate template (Figure 5) fixes endpoint structure, routing, schemas, and cross-cutting concerns such as authentication, logging, and error handling. The LLM is restricted to instantiating endpoint-specific logic within predefined slots, guided by the prompt shown in Figure 6. Validation is delegated to FastAPI’s schema-driven mechanisms, ensuring consistent enforcement of constraints specified in the OpenAPI document.
@@ 3.6. Deployed System
After iterative optimization, the final system replaces all manual translation steps with automated services and reaches a stable configuration validated by domain experts. The resulting workflow, shown in the lower-right of Figure 2, consists of three server nodes in a linear pipeline. Compared with the initial manual workflow, it achieves a net complexity reduction of 5, a communication quality improvement of 6, and an automation increase of 3.
@@
The system takes as input an initial workflow graph $\mathcal{G}^{(0)}$ with dense manual coordination and a set of domain artifacts, including OpenAPI specifications, signal definitions, and constraint documents. It outputs an optimized workflow graph $\mathcal{G}^{*}$ in which manual translation nodes are replaced by service nodes and redundant coordination edges are removed, yielding a compact executable pipeline.
@@
Each node in $\mathcal{G}^{*}$ is deployed as an independent server with a well-defined API and composed sequentially to mirror the optimized workflow graph. The architecture decouples workflow logic from the underlying language model, enabling substitution of open-source or proprietary LLMs under different cost, latency, and privacy constraints. A unified human–machine interface allows domain experts to inspect API-to-signal mappings, review generated code, and request regeneration with added constraints, preserving human oversight without manual translation.
@@
The system supports both fully automated and human-in-the-loop execution. For well-specified tasks, the pipeline runs end-to-end without intervention; ambiguous mappings or specification inconsistencies are flagged for expert review. By replacing input artifacts or domain-specific instructions, the framework can be adapted to new APIs, vehicle platforms, or other translation-heavy workflows.
@@ 4.1. Experimental Setup
The baseline implementations were reviewed and deployed in production and thus reflect accepted system behavior. In this domain, semantic errors typically appear as enum mismatches, value range violations, or missing signal bindings, all targeted by our validation and automated debugging mechanisms. We therefore adopt a conservative matching strategy, flagging ambiguous cases for human review rather than generating potentially incorrect code.
To address RQ4, we invited all practitioners with direct experience of both workflows to provide structured feedback: four domain experts and two developers. This complete set of qualified evaluators rated the system on role-specific criteria using a 5-point scale. Experts assessed communication effort, functional coverage, and implementation accuracy; developers assessed communication efficiency, debugging effort, and code maintainability.
@@
Because no existing automated solutions target API development in industrial automotive settings, we compare our system with APIs developed by human engineers assisted by proprietary AI coding agents. In this baseline, engineers were allowed to use GitHub Copilot. We assess both API quality and development time.
@@
The system separates offline processing from deployment-time execution to manage LLM usage efficiently. All compute-intensive LLM operations occur offline and do not affect runtime performance. In our experiments, we use GPT-4o (2024-05-13).
During offline processing, the LLM scans and filters the CAN database and derives read/write equations for relevant signals. In a full run involving 192 APIs and 776 CAN signals, the system made about 15,000 LLM calls over 20.6 hours, averaging 396 seconds and 80 calls per API.
During deployment, the generated read/write equations are reused. The LLM is invoked only for lightweight tasks such as matching API properties to signals and synthesizing the final endpoint. On average, each API requires about 5 LLM calls, taking roughly 10 seconds. This overhead is low enough for practical engineering workflows.
Table 3 summarizes code quality results across all six automotive domains. The automated workflow performs strongly throughout: four domains (Driver productivity, Energy, Visibility, and Dynamics) achieve perfect precision of 100%, meaning all generated property-level code was correct. The other two domains (Connected systems and Vehicle system) still achieve precision above 94%.
Recall is slightly lower, ranging from 86.5% to 95.7% across domains. This reflects our conservative property-to-signal alignment strategy, which applies strict filtering to retain only high-confidence matches. Properties with ambiguous signal mappings are flagged for human review rather than used to generate potentially incorrect code. Even so, average recall exceeds 90%, indicating broad coverage of the original specifications.
The overall F1 score of 93.7% shows that the automated workflow produces code quality comparable to baseline development across diverse automotive domains. Performance also remains stable across different expert roles: the system handles APIs requiring input from control engineers (Dynamics, Energy) as effectively as those involving HMI/UX designers or business translators.
Stakeholder satisfaction scores (Section 4.4) provide further validation: both experts and developers report high satisfaction and note that the system supports transparent monitoring through well-defined interfaces.
@@ 5. Discussion
Although Section 3 provides technical criteria for workflow automation, selecting suitable workflows ultimately still requires informed judgment. Beyond technical feasibility, successful adoption also depends on organizational readiness, stakeholder alignment, and openness to change, all of which are often difficult to evaluate in advance. In the case of spapi, the workflow was not only well-scoped but also supported by a team willing to experiment and iteratively refine the process, which proved critical to achieving tangible and sustained benefits.
@@
Our results show that LLMs can support multiple stages of workflow automation, including parsing, translation, generation, and evaluation. However, effective use of LLMs still requires experience and engineering judgment, particularly when deciding whether a task should be handled by an LLM alone or by a hybrid solution combining scripts and LLMs. The iterative nature of our approach allows LLM usage to evolve alongside the workflow itself, as teams gradually refine these design choices and improve the balance between automation, control, and reliability over time.
@@ 7. Conclusion
In this paper, we show that LLM-powered workflow automation is effective in production MSD settings. By modeling workflows as graphs and selectively automating translation steps, we substantially reduce coordination overhead while preserving human oversight throughout the process. The spapi deployment at Volvo Group validates this approach at scale. The pattern we address is not automotive-specific; it applies more broadly wherever software encodes specialized domain knowledge.